# Generalized Clausius inequalities and entanglement production in holographic two-dimensional CFTs

### v1

@@ 1 Introduction
We can thus readily motivate a systematic study of which transitions between quantum equilibrium states are physically realizable especially via a quench involving instantaneous and (in)homogeneous transfer of energy and momentum from an infinite memoryless bath. Note that such a quench is a unitary process in which the Hamiltonian changes rapidly in time. Given that quantum equilibrium states are realizable via unitary transformations of thermal states carrying momenta, they can be used to encode quantum information at finite temperature. Therefore, a study of the feasibility of such transitions between quantum equilibrium states can lead to the understanding of new principles for the construction of quantum memory and gates operating at finite temperature and in which the errors occurring over microscopic time scales can be bounded in a controlled way. As for instance, in Banerjee:2022dgv , it was argued that certain dense non-isometric encodings can be fully tolerant to fast deletion at ambient room temperature due to the quantum correction to the Landauer bound.
@@ 1.1 Summary of results
Crucially, the transitions where the total momentum density is unchanged are always allowed. We will discuss in Sec. 8 that this is consistent with the proof of the QNEC in holographic field theories. We note that the bulk energy-momentum tensor localized on the null shock is determined completely by the final and initial states. Generically, whenever the Clausius inequality holds in the dual field theory the classical null energy condition (NEC) is always satisfied by the bulk energy-momentum tensor. However, the QNEC can be violated in the dual field theory even when the energy-momentum tensor on the shock satisfies NEC. We will discuss the implications of this in Sec. 8.
@@ 2.1 Bañados-Vaidya metrics
$\langle t_{\pm\pm}\rangle=\frac{c}{12\pi}\left(\theta(-u){\mu_{\pm}^{i}}^{2}+\theta(u){\mu_{\pm}^{i}}^{2}\right)$ (18)
for transitions between homogeneous thermal states. Gravitational constraints (2.1) imply the Ward identity
@@ 4.1 Early time quadratic growth regime
where $\Delta\varepsilon$ is the change in the energy density as the result of the quench. Thus, at early times the rate of growth of entanglement is determined simply by the change in energy density in an arbitrary quench and is independent of the entangling length $l$. Our result for $\mathcal{D}$ agrees with that in Liu:2013iza ; Liu:2013qca ; Hubeny:2013hz in the limit when both the initial and final spacetimes are non-rotating BTZ black branes. Our result was perturbatively established for transitions between rotating BTZ spacetimes in Kibe:2021qjy assuming small $\mu_{\pm}^{f}-\mu_{\pm}^{i}$. The early-time quadratic growth of entanglement entropy has also been seen in numerical simulations of lattice systems Unanyan:2010 ; Unanyan:2014 .
@@ 4.2 Linear growth regime and the double scaling (steady state) limit
This result was shown analytically for transitions between non-rotating BTZ spacetimes Liu:2013iza ; Liu:2013qca ; Hubeny:2013hz . In Kibe:2021qjy , this result was also obtained in the approximation that the initial and final $\mu_{\pm}$ are close to each other. Remarkably, we find that the result is valid generally beyond the latter approximation.
@@ 6.2 The double-scaling limit
Since the double scaling limit captures very accurately the behavior of the entanglement entropy and its derivatives with respect to displacements of one of its endpoints for large intervals $\mu_{\pm}^{i,f}l\geq 1$ at intermediate times as reported before, we learn that the information of the initial state cannot be easily recovered from the latter data. As the entanglement entropy of smaller intervals thermalize quickly, we also learn that the information recovery of the initial state is difficult generally at late times. To understand more details of the scrambling of information which makes recovery harder, it should be useful to study the Renyi entropies in the double scaling limit and their derivatives under displacements of one of the endpoints.
@@ 8 Discussion
The proof of the QNEC in holographic field theories in Koeller_2016 shows that if the NEC is satisfied in the bulk of a smooth classical geometry which is the solution of a two-derivative gravitational theory, then it should imply that the QNEC is satisfied in the dual field theory. This however does not imply anything for a solution with a null shock unless it is the limit of a smooth solution of a two-derivative gravity theory. When momentum density vanishes at all times, the Bañados-Vaidya geometry with the null shock is indeed a limit of a smooth solution of a massless scalar field coupled to gravity as discussed in Sec. 2. Therefore, our finding that the QNEC is satisfied when the change in momentum density vanishes, is consistent with the proof. Our general results suggest that the Bañados-Vaidya geometries where the NEC is satisfied in the bulk but the QNEC is violated in the dual field theory cannot be realized as limits of smooth solutions of a two-derivative gravity theory, i.e. Einstein’s gravity minimally coupled to matter fields. Therefore, such transitions should not be possible in the large $N$ and strong coupling limit in which the dual gravitational theory is Einstein’s gravity minimally coupled to matter fields, even if we allow the process to be realized in finite time (otherwise we should be able to take a delta function limit for the source at the boundary). However, the disallowed transition may be possible in the full classical string field theory which is the dual gravitational description of the holographic CFT at finite coupling and large $N$. In the latter case, the HRT prescription itself should be modified.
@@
Pertinently, it has been shown in Banerjee:2024sqq that one can obtain the Nambu-Goto equation from 2+1-D gravity via the junction conditions. [^10] This result may not generalize to higher dimensions. As shown in Banerjee:2024sqq , the independent equations for the junction conditions outnumber the degrees of freedom of the junction in higher dimensions. So, the worldvolume extremization equations for the co-dimension one brane constituting the junction may not follow from the junction conditions. More precisely, the general solutions of 2+1-D gravity with a junction have one-to-one correspondence with the solutions of the non-linear Nambu-Goto equations up to a finite number of rigid parameters related to worldsheet and spacetime isometries. These solutions of 2+1-D gravity are dual to novel types of interfaces in the universal sector of 1+1-D CFTs featuring time-reparametrization at the interface. (In fact, one can decode the bulk Nambu-Goto solution from the time-reparametrization at the interface Banerjee:2024sqq ). These solutions give a method for constructing a large variety of quantum processors (giving a map between the Hilbert spaces of the CFT on the left and right sides of the interface) utilizing the fundamental string junction in 2+1-D gravity.
@@
Such solutions of 2+1-D gravity with the fundamental string junction, however, are not limits of smooth solutions of theories where bulk matter fields (with the matter action involving only two derivatives) are coupled minimally to Einstein’s gravity. Therefore, we can expect that our methods when extended to such solutions will produce generalized Clausius inequalities via the validity of QNEC constraining quantum processors which can be physically realized via the interfaces dual to fundamental string junctions in 2+1-D gravity. Such results should give us new fundamental insights into quantum gravity. This can be expected as the QNEC itself was conjectured as a way to satisfy the generalized second law in semi-classical gravity Bousso_2016 .
@@
Another important result of our work is that the initial state can be recovered from the entanglement entropy of different spatial intervals (which are on a constant time slice in the frame of observers who have finite boosts with respect to the final equilibrium state) at early times after the quench but not at late times particularly in the regime where the entanglement entropy thermalizes at the speed of light. This leads us to pose an interesting question about the scrambling of quantum information in holographic CFTs. Specifically, we can ask if the (time dependent) entanglement spectrum of an interval of a fixed finite length $l$ can be used to recover the initial state at arbitrary times after the quench which is fundamentally a unitary process. It would be of interest to study the time-dependent Renyi entropies of intervals of fixed lengths during quenches leading to transitions between thermal states with uniform momentum density especially in holographic theories, and examine if the initial state can be recovered from the entanglement spectrum at an arbitrary time after the quench. This should give us a way to precisely quantify how hard it is to recover the information of the initial state with the passage of time, i.e. how fast information scrambles and also the associated length scales of scrambling.
@@ Acknowledgments
We would like to thank Costas Bachas, Giuseppe Policastro, Mukund Rangamani and Ronak Soni for valuable discussions. The research of TK and PR is supported by NRF Postdoctoral Fellowships. The research of AM has been supported by Fondecyt grant 1240955.

### v2

# Wits, 2050, South Africa

@@ 1 Introduction
We can thus readily motivate a systematic study of which transitions between quantum equilibrium states are physically realizable especially via a quench involving instantaneous and (in)homogeneous transfer of energy and momentum from an infinite memoryless bath. [^5] When a system is in contact with an infinite bath, it is expected to attain thermal equilibrium without changing the bath. When the bath is also memomoryless, the auto-correlation functions of the bath vanish, and the system is expected to undergo unitary evolution with a time-dependent Hamiltonian. Note that for an isolated system, in contrast, the Hamiltonian is independent of time. Note that such a quench is a unitary process in which the Hamiltonian changes rapidly in time. Given that quantum equilibrium states are realizable via unitary transformations of thermal states carrying momenta, they can be used to encode quantum information at finite temperature. Therefore, a study of the feasibility of such transitions between quantum equilibrium states can lead to the understanding of new principles for the construction of quantum memory and gates operating at finite temperature and in which the errors occurring over microscopic time scales can be bounded in a controlled way. As for instance, in Banerjee:2022dgv , it was argued that certain dense non-isometric encodings can be fully tolerant to fast deletion at ambient room temperature due to the quantum correction to the Landauer bound.
@@ 1.1 Summary of results
Crucially, the transitions where the total momentum density is unchanged are always allowed. We will discuss in Sec. 8 that this is consistent with the proof of the QNEC in holographic field theories. We note that the bulk energy-momentum tensor localized on the null shock is determined completely by the final and initial states. Generically, whenever the Clausius inequality holds in the dual field theory the classical null energy condition (NEC) is always satisfied by the bulk energy-momentum tensor. However, the QNEC can be violated in the dual field theory even when the energy-momentum tensor on the shock satisfies NEC (and the bulk also has no naked singularities, c.f. Sec. 2). We will discuss the implications of this in Sec. 8.
@@ 2.1 Bañados-Vaidya metrics
$\langle t_{\pm\pm}\rangle=\frac{c}{12\pi}\left(\theta(-u){\mu_{\pm}^{i}}^{2}+\theta(u){\mu_{\pm}^{f}}^{2}\right)$ (18)
for transitions between homogeneous thermal states. It follows that the (homogeneous) energy density $\epsilon$ is given by
$\epsilon=\langle t_{tt}\rangle=\frac{c}{12\pi}\left(\theta(-u)\left({\mu_{+}^{i}}^{2}+{\mu_{-}^{i}}^{2}\right)+\theta(u)\left({\mu_{+}^{f}}^{2}+{\mu_{-}^{f}}^{2}\right)\right)$ (19)
Gravitational constraints (2.1) imply the Ward identity
@@ 4.1 Early time quadratic growth regime
where $\Delta\varepsilon$ is the change in the energy density as the result of the quench (c.f. Eq. (19)). Thus, at early times the rate of growth of entanglement is determined simply by the change in energy density in an arbitrary quench and is independent of the entangling length $l$. Our result for $\mathcal{D}$ agrees with that in Liu:2013iza ; Liu:2013qca ; Hubeny:2013hz ; Ageev:2017wet in the limit when both the initial and final spacetimes are non-rotating BTZ black branes. Our result was perturbatively established for transitions between rotating BTZ spacetimes in Kibe:2021qjy assuming small $\mu_{\pm}^{f}-\mu_{\pm}^{i}$. The early-time quadratic growth of entanglement entropy has also been seen in numerical simulations of lattice systems Unanyan:2010 ; Unanyan:2014 .
@@ 4.2 Linear growth regime and the double scaling (steady state) limit
This result was shown analytically for transitions between non-rotating BTZ spacetimes Liu:2013iza ; Liu:2013qca ; Hubeny:2013hz ; Ageev:2017wet . In Kibe:2021qjy , this result was also obtained in the approximation that the initial and final $\mu_{\pm}$ are close to each other. Remarkably, we find that the result is valid generally beyond the latter approximation.
@@ 6.2 The double-scaling limit
Since the double scaling limit captures very accurately the behavior of the entanglement entropy and its derivatives with respect to displacements of one of its endpoints for large intervals $\mu_{\pm}^{i,f}l\geq 1$ at intermediate times as reported before, we learn that the information of the initial state cannot be easily recovered from the latter data. As the entanglement entropy of smaller intervals thermalize quickly, we also learn that the information recovery of the initial state is difficult generally at late times. To understand more details of the scrambling of information which makes recovery harder, it should be useful to study the Renyi entropies in the double scaling limit and their derivatives under displacements of one of the endpoints (since these have more fine-grained information than the entanglement entropy).
@@ 8 Discussion
The proof of the QNEC in holographic field theories in Koeller_2016 shows that if the NEC is satisfied in the bulk of a smooth classical geometry which is the solution of a two-derivative gravitational theory, then it should imply that the QNEC is satisfied in the dual field theory. However this does not imply anything for a solution with a null shock unless it is the limit of a smooth solution of a two-derivative gravity theory. When momentum density vanishes at all times, the BTZ-Vaidya geometry with the null shock is in fact a limit of a smooth solution of a massless scalar field coupled to gravity as discussed in Sec. 2. Therefore, our finding that the QNEC is satisfied when the change in momentum density vanishes, is consistent with the proof. Our general results suggest that the Bañados-Vaidya geometries where the NEC is satisfied in the bulk but the QNEC is violated in the dual field theory cannot be realized as limits of smooth solutions of a two-derivative gravity theory, i.e. Einstein’s gravity minimally coupled to matter fields. Therefore, even if such transitions are possible in finite time at large $N$ and strong coupling (so that the dual gravitational theory is Einstein’s gravity minimally coupled to matter fields), it should not be possible to take the limit in which the transition becomes instantaneous in the bulk. [^11] Consider sourcing the dual $1+1$-D CFT with a gravitational potential via $g_{tt}=-1-e^{-2\phi(x)}f(t)$ with $f(t)$ being a Gaussian function. This leads to a gravitational force acting for finite time leading to the injection of net momentum to the thermal matter. However, even in the limit in which $f(t)$ becomes a delta function, the transition of the system to the final equilibrium with the added uniform momentum density can occur over a period of time. It should be possible to study this explicitly via generalized hydrodynamics Doyon:2023zcl which can explain how the right moving excitations attain different momentum from the left movers. It would also be interesting to study this in a holographic setup. Furthermore, we should study examples in which the modification of the boundary metric is not required and understand generally if the transition can be made instantaneously in an appropriate limit when it is not forbidden at large $N$ and strong coupling. We thank the referee for a relevant discussion. However, the disallowed transition with an instantaneous limit in the bulk may be possible in the full classical string field theory which is the dual gravitational description of the holographic CFT at finite coupling and large $N$. In the latter case, the HRT prescription itself should be modified.
@@
Pertinently, it has been shown in Banerjee:2024sqq that one can obtain the Nambu-Goto equation from 2+1-D gravity via the junction conditions. [^12] This result may not generalize to higher dimensions. As shown in Banerjee:2024sqq , the independent equations for the junction conditions outnumber the degrees of freedom of the junction in higher dimensions ($D>3$). So, although some special symmetric solutions of the world-volume extremization equations can be expected to be realized in the limit where the brane tension vanishes, we should not be able to obtain a generic solution of these equations in the tensionless limit for $D>3$. More precisely, the general solutions of 2+1-D gravity with a junction have one-to-one correspondence with the solutions of the non-linear Nambu-Goto equations up to a finite number of rigid parameters related to worldsheet and spacetime isometries. These solutions of 2+1-D gravity are dual to novel types of interfaces in the universal sector of 1+1-D CFTs featuring time-reparametrization at the interface. (In fact, one can decode the bulk Nambu-Goto solution from the time-reparametrization at the interface Banerjee:2024sqq ). These solutions give a method for constructing a large variety of quantum processors (giving a map between the Hilbert spaces of the CFT on the left and right sides of the interface) utilizing the fundamental string junction in 2+1-D gravity.
@@
Such solutions of 2+1-D gravity with the fundamental string junction, however, are not expected to be limits of smooth solutions of theories where bulk matter fields (with the matter action involving only two derivatives) are coupled minimally to Einstein’s gravity. Therefore, we can expect that our methods when extended to such solutions will produce generalized Clausius inequalities via the validity of QNEC constraining quantum processors which can be physically realized via the interfaces dual to fundamental string junctions in 2+1-D gravity. Such results should give us new fundamental insights into quantum gravity. This can be expected as the QNEC itself was conjectured as a way to satisfy the generalized second law in semi-classical gravity Bousso_2016 .
@@
Another important result of our work is that the initial state can be recovered from the entanglement entropy of different spatial intervals (which are on a constant time slice in the frame of observers who have finite boosts with respect to the final equilibrium state) at early times after the quench but not at late times particularly in the regime where the entanglement entropy thermalizes at the speed of light. This leads us to pose an interesting question about the scrambling of quantum information in holographic CFTs. Specifically, we can ask if the (time dependent) entanglement spectrum of various (space-like) intervals of a fixed finite length $l$ can be used to recover the initial state at arbitrary times after the quench which is fundamentally a unitary process. It would be interesting to study the time evolution of the Renyi entropies for this purpose as these have more fine grained information about the entanglement spectrum. This should give us a way to precisely quantify information scrambling in a subsystem. Furthermore, we should also be able to extract how the time-scale of scrambling of information in an interval (subsystem) depends on its length (size).
@@
A variety of explicit quench protocols have been studied in CFTs (see Bernard:2016nci and references therein). A particularly interesting quench protocol involves instantaneously joining two half-line CFTs prepared at different temperatures leading to the formation of a steady state Bernard:2016nci . This setup has also been studied holographically in Erdmenger:2017gdk . It would be interesting to study if QNEC imposes any bounds on such quenches leading to the formation of steady states.
Finally, from the holographic point of view, it should be interesting to understand the necessary and sufficient conditions (beyond the classical NEC) which the stress tensor localized on a bulk shock or junction should satisfy generally so that the QNEC is also satisfied in the corresponding process in the dual field theory. Such a set of conditions for the bulk stress tensor on the co-dimension one hypersurface can be non-local.
@@ Acknowledgments
We would like to thank Costas Bachas, Giuseppe Policastro, Mukund Rangamani and Ronak Soni for valuable discussions. The research of PR is supported by the NRF Postdoctoral Fellowship. The research of AM has been supported by Fondecyt grant 1240955.